OpenStudy (anonymous):

-2n-2 ___________=12 3 What is n?

sammixboo (sammixboo):

Does your problem look like this? \(\rm \color{red}{\dfrac{-2n~-~2}{3}=~12}\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@sammixboo yes

OpenStudy (ageta):

et's solve your equation step-by-step. −2n−2/3=12 Step 1: Simplify both sides of the equation. −2/3 n+ −2/3 =12 Step 2: Add 2/3 to both sides. −2/3n+ −2/3+2/3 =12+2/3 −2/3 n=38/3 Step 3: Divide both sides by (-2)/3. −2/3n −2/3=38/3 −2/3 n=−19

OpenStudy (anonymous):

@ageta thank you so much! I forgot to do step 1

OpenStudy (ageta):

use this http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

@ageta thanks i bookmarked it!

OpenStudy (ageta):

good

OpenStudy (anonymous):

use this tool :- http://www.acalculator.com/math-calculators.html